OpenStudy (anonymous):

With Quadratic Functions how do you go from Standard Form to Intercept Form?

OpenStudy (anonymous):

Intercept form: f(x) = a(x – p)(x – q). The points x = p and x = q are the x-intercepts. (factored form)

OpenStudy (anonymous):

I know intercept form but idk how to go from y=x^2-8x+12(standard form) to y=a*(x-r1)(x-r2)(intercept form)

OpenStudy (amistre64):

intercept form is just the factors ... you can find the roots of an equation using a variety of methods if it is not easily factored.